OpenStudy (allieatthedisco):

if x varies jointly as y and z, and x=10 when y=4 and z=15, find x when y=8 and z=20

OpenStudy (p0sitr0n):

do you have more points? because it could be many things from that setup x could be 13 y: 4->8 = +4 z: 15->20 +5 so x: 10->13 +3 just a guess

OpenStudy (allieatthedisco):

I'm afraid those are the only points i was given

OpenStudy (nikato):

So x varies jointly as y and z would be x=k(yz) Plugging in x=10, y=4, z=15 10=k(4)(15) Now solve for k

OpenStudy (allieatthedisco):

so k would = .16?

OpenStudy (nikato):

Not quite. k=1/6 10=60k. Divide 60 So k=1/6

OpenStudy (nikato):

So you have x=1/6 (yz) And you have y=8 and z=20 And you're trying to solve x. So plug the numbers in x=1/6 (8)(20) So what's x?